If on an average 1 ship in every 15 sinks, find the chance that out of 10 ships atleast 6 will arrive safely.

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
given q=1/15 so p=1-q=14/15
so distribution is (14/15+1/15)10
Atleast 6 will arrive safely=p(x=6)+p(x=7)+p(x=8)+p(x=9)+p(x=10)
p(x=6)=10c6(14/15)6(1/15)4
similarly by adding the all the terms you will get the answer
Thanks and Regards
Shaik Aasif
askIITians faculty